# Path Loss Prediction Using Deep Learning

### v2

@@ Abstract
Radio deployments and spectrum planning benefit from path loss predictions. Obstructions along a communications link are often considered implicitly or through derived metrics such as representative clutter height or total obstruction depth. In this paper, we propose a path-specific path loss prediction method that uses convolutional neural networks to automatically perform feature extraction from high-resolution obstruction height maps. Our methods result in low prediction error in a variety of environments without requiring derived metrics.
@@ I Introduction
Path loss modeling is a tool used by the wireless industry and regulatory bodies to optimize wireless coverage and avoid unwanted interference. Path loss models predict point-to-point (P2P) wireless signal attenuation along the communications link from the transmitter (Tx) to the receiver (Rx). Path loss is primarily affected by frequency, link distance, and obstructions in the line of sight (LOS). In LOS conditions, free space path loss (FSPL), which uses frequency and distance, is sufficient for accurate P2P path loss predictions. In non-LOS conditions, however, FSPL typically underestimates path loss.
@@
Previous work in [3] shows improvement over P.1812-6 in non-LOS conditions, by capturing GIS information in one novel feature: total obstruction depth. This scalar feature is computed using surface information along the direct path. Alongside frequency and distance, obstruction depth yields a root mean square error (RMSE) in the range of 6-8 dB as compared to 8-13 dB for P.1812-6 [3]. While using total obstruction depth as a feature can produce lower errors, it does not explicitly consider the surface height profile, which can influence path loss in more complicated ways due to diffraction and Fresnel zone effects [4].
The goal of this paper is to explicitly utilize the surface height profile to train a path-specific propagation model. Leveraging high-resolution clutter maps and deep neural networks to directly predict path loss, we aim to study whether high-dimensional spatial features can be extracted more reliably by neural networks than by human-defined feature engineering.
@@
This paper describes path loss models trained on 2-D surface height data alongside measurement data. Section II describes the pre-processed datasets used to train and evaluate the path loss models. Section III explains the multi-dimensional features extracted from the data and the model architecture. Section IV describes the training and evaluation process and results. Finally, Section V summarizes our findings.
@@ III-B Model Architecture
Given the spatial structure of the data, we employ convolutional neural networks (CNNs) [16], treating the data as images. Each convolution layer has a kernel size of 3$\times$3, stride size of 2$\times$2, and zero-padding size of 1$\times$1. Each max pool layer has a kernel size of 3$\times$3, stride size of 1$\times$1, and “same” zero-padding [16].
During training, we use dropout at a rate of 25% on the 4096 flattened features to minimize overfitting. The extracted features are input to a fully connected network (FCN) [16] with 256 hidden units and rectified linear unit (ReLU) activation. The final output is path loss. Given the low RMSE in [3] with only 3 features, 4096 features are likely to be sufficient, regardless of the feature extraction process.
@@ IV-B Evaluation of Result Consistency
Table III summarizes the cross-validation results with the mean and standard deviation (SD) of test RMSE for each holdout. We also compare the mean results with ITU-R P.1812-6 [2] (the industry standard for 1-6 GHz, hereafter referred to as P.1812) and a three-feature FCN using frequency, distance, and total obstruction depth [3].
@@
The model generally produces lower RMSE for lower frequencies and shorter distances, as shown in Table IV. The major errors occur on the extremely long links. Across all runs and holdouts, the CNN achieves 7.37 dB RMSE on the shortest 99.9% of unique holdout links ($<$45 km), and 12.04 dB RMSE on the remaining 0.1% ($<$800 links). This implies that re-sampling has a significant effect on performance, especially for longer links. Future work should ensure more representation of long links in the training set, or only operate the model on shorter links to ensure confidence in prediction.
TABLE IV: Subset RMSEs Computed Across All Runs
Subset Holdout RMSE [dB]
449 MHz 6.31
915 MHz 7.10
1802 MHz 7.24
2695 MHz 7.32
3602 MHz 7.99
5850 MHz 8.15
*Shortest 99.9% 7.37
Longest 0.1% 12.04
*99.9% of links are below 45 km.
Compared to the FCN, our model shows lower mean RMSE on four of six holdouts, with higher mean RMSE on the other two holdouts (London and Merthyr Tydfil). Although four of six holdouts show lower RMSE than the FCN, the fact that other holdouts have higher RMSE in the presence of richer feature sets suggests a further need for model refinement and feature selection. Though these trends may differ when attempting to generalize to other countries and measurements, using multi-city cross-validation suggests that the result would likely generalize to other unseen regions.
Our methods simplify modeling by eliminating feature engineering but require a larger initial investment in training time–80 minutes per run vs. 2 minutes for the FCN. The inference speed of the CNN and the FCN are 4000 and 10 000 samples/s, respectively, with the same hardware (Amazon Web Services g5.4xlarge), which are both computationally feasible.
Our CNN-based approach provides significant improvement over the current industry standard, achieving lower RMSE than P.1812 in all six holdouts. This implies that higher-dimensional information outperforms representative clutter heights in predictive models.
@@ V Conclusion
Path loss modeling is an asset for effective radio deployments and spectrum planning. This paper described the use of CNNs in conjunction with surface path profiles to train a path-specific path loss model with low test RMSE in a variety of environments. We showed that CNNs trained on 2-D path profiles outperform a modern FCN path loss model with equivalent scalar features in four of six holdouts. These CNNs also consistently achieved lower RMSE than the industry standard in a variety of communications environments, without the need for derived clutter metrics or scalar features.

### v3

# Map-Based Path Loss Prediction in Multiple Cities Using Convolutional Neural Networks

@@ Abstract
Radio deployments and spectrum planning benefit from path loss predictions. Obstructions along a communications link are often considered implicitly or through derived metrics such as representative clutter height or total obstruction depth. In this paper, we propose a path-specific path loss prediction method that uses convolutional neural networks to automatically perform feature extraction from 2-D obstruction height maps. Our methods result in low prediction error in a variety of environments without requiring derived metrics.
@@ I Introduction
Path loss modeling is a tool used by the wireless industry and regulatory bodies to optimize wireless coverage and avoid unwanted interference. Path loss models predict point-to-point (P2P) wireless signal attenuation along a communications link from the transmitter (Tx) to the receiver (Rx). Path loss is primarily affected by frequency, link distance, and obstructions in the line of sight (LOS). In LOS conditions, free space path loss (FSPL), which uses frequency and distance, is sufficient for accurate P2P path loss predictions. In non-LOS conditions, however, FSPL typically underestimates path loss.
@@
Previous work in [3] shows improvement over P.1812-6 in non-LOS conditions, by capturing GIS information in one novel feature: total obstruction depth. This scalar feature is computed using surface information along the direct path. Alongside frequency and distance, obstruction depth yields a root mean square error (RMSE) in the range of 6-8 dB as compared to 8-13 dB for P.1812-6 [3]. While using total obstruction depth as a feature can produce low errors, it does not explicitly consider the surface height profile, which can influence path loss in more complicated ways due to diffraction and Fresnel zone effects [4]. Designing and computing engineered features can also be time-consuming or difficult.
The goal of this paper is to explicitly utilize the surface height profile to train a path-specific propagation model akin to P.1812 (unlike a model trained only on derived features [3]). Rather than aiming to reduce RMSE compared to [3], our objective is to study whether deep neural networks can automatically extract spatial features from GIS data to streamline the process over feature engineering. Leveraging high-resolution clutter maps and deep learning to directly predict path loss, this paper will determine whether automatic feature extraction and regression integrated in a single neural network can achieve RMSE lower than P.1812 and similar to [3].
@@
This paper describes path loss models trained on 2-D obstruction height maps alongside measurement data. Section II describes the preparation of training and evaluation data. Section III describes the 2-D features and the model architecture. Section IV describes training and evaluation, and Section V summarizes our findings.
@@ III-B Model Architecture
Given the spatial structure of the data, we employ convolutional neural networks (CNNs) [16], treating the data as images. Each convolution layer has a kernel size of 3$\times$3, stride size of 2$\times$2, and zero-padding size of 1$\times$1. Each max pool layer has a kernel size of 3$\times$3, stride size of 1$\times$1, and “same” zero-padding [16]. During training, we use dropout at a rate of 25% on the 4096 flattened features to minimize overfitting. The extracted features are input to a fully connected network (FCN) [16] with 256 hidden units and rectified linear unit (ReLU) activation. The final output is path loss. Given the low RMSE in [3] with only 3 features, 4096 features are likely to be sufficient, regardless of the feature extraction process.
@@ IV-B Evaluation of Result Consistency
Table III summarizes the cross-validation results with the mean and standard deviation (SD) of test RMSE for each holdout. We also compare the means with ITU-R P.1812-6 [2] (the industry standard for 1-6 GHz) and a three-feature FCN using frequency, distance, and total obstruction depth [3].
@@
The model RMSE increases monotonically with frequency, showing 6.31 and 8.15 dB on 449 and 5850 MHz respectively. RMSE also increases with distance, with the largest errors occurring on the extremely long links. Across all runs and holdouts, the CNN achieves 7.37 dB RMSE on the shortest 99.9% of unique holdout links ($<$45 km), and 12.04 dB RMSE on the remaining 0.1% ($<$800 links). This implies that re-sampling can have a significant effect on performance, especially for longer links. Future work should ensure more representation of long links in the training set, or only operate the model on shorter links to ensure confidence in prediction.
Compared to the FCN, our CNN shows lower mean RMSE on four of six holdouts, with higher mean RMSE on London and Merthyr Tydfil. Though these trends may differ when attempting to generalize to other countries and measurements, using multi-city cross-validation suggests that the result would likely generalize to other unseen regions. Neither model provides strictly lower RMSE, with similar aggregate results. The choice of model depends on the user’s priorities: The CNN provides the novelty of streamlining modeling by eliminating feature engineering but requires a larger initial investment in training time–80 minutes per run vs. 2 minutes for the FCN. The inference speed of the CNN and the FCN are 4000 and 10 000 samples/s, respectively, with the same hardware (Amazon Web Services g5.4xlarge), which are both practical speeds. Finally, the fact that distinct modeling approaches applied to the same DSM achieve similarly low RMSE validates the use of DSM in machine learning-based path loss models, whether features are extracted automatically (CNN) or manually (FCN).
The CNN showing similar results to the FCN with richer inputs suggests room for further model refinement to reduce overfitting (as is often the case in machine learning research). Still, our novel CNN-based approach provides significant improvement over the current industry standard, achieving lower RMSE than P.1812 in all six holdouts. This implies that higher-dimensional information outperforms representative clutter heights in predictive models.
@@ V Conclusion
Path loss modeling is an asset for effective radio deployments and spectrum planning. This paper described the use of CNNs in conjunction with surface path profiles to train a path-specific path loss model with low test RMSE in a variety of environments. We showed that CNNs trained to provide automatic feature extraction from 2-D path profiles achieve similar RMSE to a modern FCN path loss model with scalar features. These CNNs consistently achieved lower RMSE than the industry standard in a variety of communications environments, without the need for derived clutter metrics or scalar GIS features.